OpenStudy (anonymous):

At the school cafeteria, the cost of one milk is $0.50. The total cost of 5 turkey sandwiches and 4 milks is the same as the total cost of 4 turkey sandwiches and 8 milks.

OpenStudy (anonymous):

are you trying to find the cost of turkey sandwhich?

OpenStudy (anonymous):

Sorry for the missing information, parent trying to re-learn to help my son. I need to: a. Write an equation that can be used to find the cost of a turkey sandwich. Explain what each expression or value represents in the equation. b. Use the equation you wrote in part a. to find the cost of a turkey sandwich. Explain how you solved the equation. I *think* the equation (inequality) should be: T= turkey sandwich 5t + 4(0.50) = 4t + 8(0.50) 5t + 2.00 = 4t + 4.00 Is that right?

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

bingoo

OpenStudy (anonymous):

Yay! This stuff is not easy, I have been out of school for 30 years :/

OpenStudy (anonymous):

practice makes perfect :D

OpenStudy (zzr0ck3r):

It was no easier 30 years ago:P

OpenStudy (anonymous):

No, it wasn't. Math was never my subject. Trying to solve this now, but not sure of how to. So let me run this by you... 5t+2-4t=4t+4-4t ??

OpenStudy (anonymous):

subtract 4t from both sides then add the 2. gotta believe in yourself, you already know how to do it!

OpenStudy (anonymous):

subtract the 2*

OpenStudy (anonymous):

So, t+2 = 0 +4 -2 -2 t = 2

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

exactly. check your work by plugging it back in. 5t+4(.5) = 4t+8(.5) if t = 2 then they should be the same.. 5(2)+2 = 4(2)+4 10=10

OpenStudy (anonymous):

Thank you all very much! I probably do know how to do this stuff but I needed a little moral support. I appreciate your time :)